OpenStudy (anonymous):

Find the inverse function of f. f(x) = 7x + 2

OpenStudy (slaaibak):

y = 7x + 2 switching y and x: x = 7y + 2 In terms of y: 7y = x-2 y = (x-2)/7

OpenStudy (slaaibak):

\[f^{-1}(x) = {x-2 \over 7}\]